# Towards Robust and Realible Multimodal Fake News Detection with Incomplete Modality

### v1

@@ 4.6. Ablation Study
1) To validate the effectiveness and robustness of the modality-specfic multi-expert reasoning paradigm in MMLNet for the MFDN task, we conducted ablation studies by sequentially removing the textual expert ${L}_{c}^{h}$, the visual expert ${L}_{c}^{r}$, and the multimodal expert ${L}_{c}^{f}$. The framework achieved the best performance when all experts collaborated, demonstrating that all modal experts play indispensable roles in discerning the authenticity of news. Notably, removing the ${L}_{c}^{f}$ resulted in a significant performance drop, which highlights the critical importance of modeling the distribution of cross-modal interaction features for the success of MFND.
@@
4.7. Analysis on Modality Missing Learning.
Answear to RQ4. As shown in Figure 7(a), our modality missing learning is more suitable for MFND tasks with incomplete modalities. On the Weibo dataset with 50% text and 50% image modality missing, we conducted the following ablation experiments on modality missing learning:
@@
  • w MML, w/o Weight. MMLNet with modality missing learning but without the weight in Eq18.
@@
We visualized the multimodal features before classification via T-SNE [^5] https://github.com/mxl1990/tsne-pytorch in Figure 7. The ”w/ MML” line shows that MMLNet with modality missing learning achieves up to an 4% higher accuracy than without it, indicating our modality missing learning is beneficial in modality missing scenarios. ”w MML, w/o Weight” is second-best, showing that positive and negative samples based on news authenticity labels improve the MMLNet model’s ability to detect fake news. ”w/o MML, w/ MCL” has similar performance to ”w/o MML”, but Figure 7(d) shows that vanilla multimodal contrastive learning has the worst visualization results, as modality missing undermines vanilla multimodal contrastive learning. This occurs because when image and text modality information is missing, the vanilla multimodal contrastive learning is misled by the missing information and cannot effectively distinguish between positive and negative samples.
4.8. Case Study
@@
Answear to RQ5. In Figure 8, we provide a case study on MFND to provide a detailed analysis with some interpretability. Analyzing the left 4 columns, although most baseline models are able to distinguish fake news when the modality information is complete, they cannot make accurate judgments when the modality information is incomplete. However, MMLNet effectively corrects the impact of missing modality information on model performance by combining distributions from different modality experts. For example, when text modality absence predominates, the lack of text modality information causes erroneous judgments by the text expert, but the image modality expert and multimodal expert distributions compensate for this, improving MMLNet’s robustness in MFND.
@@
Answear to RQ6. Although MMLNet has achieved significant improvements in the incomplete modality MFND task, it still has limitations. When modality information is severely missing, it cannot effectively distinguish the authenticity of news. As shown in the sixth column of Figure 8, when the text modallity information is missing to a large extent, the model will mistakenly judge the news as true, mainly because of the loss of key person information. Similarly, when the image modal information is missing to a large extent, effective information cannot be obtained from the image modality. As shown in the eighth column of Figure 8, when the text information on the image is lost, resulting in semantic inconsistency between the image modality and the text modality, the model mistakenly judges the sample as fake news.
5. Conclusion and Discussion
@@
In this paper, we focus on the impact of modality missing information on the robustness of multimodal fake news detection. These issues are common in practical applications, yet they have been insufficiently explored in previous studies. We propose MMLNet, which models the distribution of different modalities through multi-expert learning. The modality missing adapter compensates for missing modality information, while incomplete modality learning enhances the robustness of feature distributions. Multi-expert collaboration enables accurate identification of misinformation. Extensive experiments and ablation studies under both complete and incomplete modalities demonstrate the effectiveness of our approach. Our future work will explore more missing modality scenarios in MFND tasks, such as video and audio modalities, to further improve the adaptability of our method. Meanwhile, we will study how to design more modality missingness to better fit real-world contexts.
@@ Acknowledgements.
This work was supported by the National Science Fund of China under Grant No. 62376121 and No. 62501644.
Appendix A Detailed experimental results
Table 3. Experimental results of our proposed MMLNet
Missing Rate (%) Modality Weibo Weibo21 Pheme
Text Image ACC M-F1 AUC ACC M-F1 AUC ACC M-F1 AUC
↓100 - 100 91.19 91.18 96.68 90.40 90.40 95.98 83.82 79.27 90.91
25 75 90.23 90.23 96.20 90.08 90.08 95.94 82.83 79.03 86.70
50 50 91.60 91.58 96.87 89.91 89.91 96.41 82.01 76.01 74.16
75 25 92.55 92.55 96.99 90.56 90.56 96.11 80.19 74.82 84.99
100 - 92.21 92.20 97.67 89.59 89.58 95.51 80.03 73.37 85.18
↓75 - 75 92.08 92.07 97.18 91.70 91.70 97.11 82.50 76.72 84.98
25 50 91.74 91.72 97.64 92.19 92.19 97.44 83.99 80.61 87.43
50 25 93.37 93.37 97.89 91.86 91.86 96.90 80.69 76.29 86.52
75 - 93.58 93.57 97.71 91.05 91.05 96.02 81.02 77.01 86.32
↓50 - 50 92.96 92.96 97.92 91.70 91.70 98.03 83.49 79.48 88.42
25 25 93.92 93.92 98.64 93.33 93.33 97.69 82.17 76.39 88.18
50 - 94.40 94.39 98.50 92.19 92.19 97.50 82.50 77.76 87.60
↓25 - 25 94.60 94.60 98.76 93.98 93.98 98.08 83.00 77.53 90.39
25 - 94.88 94.87 98.85 93.65 93.65 97.91 84.81 79.89 88.91
- - - 95.22 95.21 98.93 94.95 94.95 98.23 87.78 84.35 90.97
Table 4. Experimental results of NSLM
Missing Rate (%) Modality Weibo Weibo21 Pheme
Text Image ACC M-F1 AUC ACC M-F1 AUC ACC M-F1 AUC
↓100 - 100 88.73 88.69 89.03 89.26 89.26 89.26 78.87 75.86 77.97
25 75 86.07 85.83 85.74 84.87 84.72 84.96 82.50 80.01 82.41
50 50 82.59 82.25 82.23 84.06 83.98 84.01 80.52 75.70 75.18
75 25 71.74 70.52 71.16 77.39 77.39 77.39 74.09 60.08 59.69
100 - 48.94 48.28 49.35 53.00 45.82 53.30 71.28 41.61 50.00
↓75 - 75 91.33 91.32 91.34 90.89 90.89 90.89 83.99 80.06 79.50
25 50 91.39 91.39 91.46 89.26 89.28 89.28 80.52 75.79 75.35
50 25 87.16 87.11 87.04 86.34 86.26 86.40 77.88 67.37 65.61
75 - 78.56 78.52 78.51 75.12 74.71 75.23 75.57 60.55 60.21
↓50 - 50 91.67 91.67 91.73 87.64 87.58 87.59 81.18 75.89 74.79
25 25 91.19 91.18 91.15 89.59 89.57 89.63 81.84 76.83 75.77
50 - 87.03 87.01 87.01 86.66 86.63 86.71 80.03 73.37 71.75
↓25 - 25 92.42 92.42 92.44 90.08 90.05 90.04 82.50 77.31 75.89
25 - 91.80 91.80 91.82 90.73 90.73 90.73 81.84 78.73 80.06
- - - 92.28 92.28 92.32 90.89 90.88 90.92 84.65 81.82 82.88
Table 5. Experimental results of MIMoE
Missing Rate (%) Modality Weibo Weibo21 Pheme
Text Image ACC M-F1 AUC ACC M-F1 AUC ACC M-F1 AUC
↓100 - 100 90.44 90.39 90.31 88.13 88.07 88.19 80.36 72.80 87.19
25 75 90.85 90.84 91.04 89.26 89.26 89.25 77.88 67.90 83.21
50 50 84.36 84.23 84.74 86.66 86.64 86.70 73.26 58.20 79.17
75 25 80.06 80.06 80.10 82.11 82.10 82.10 74.25 61.10 75.09
100 - 74.19 73.10 73.62 79.83 79.76 79.79 72.44 55.90 72.79
↓75 - 75 91.26 91.23 91.16 90.40 90.38 90.45 80.19 74.70 85.99
25 50 91.12 91.10 91.07 91.22 91.21 91.24 80.69 73.60 85.36
50 25 89.35 89.33 89.31 88.94 88.92 88.98 78.54 72.50 80.97
75 - 80.75 80.56 80.51 79.02 78.87 79.09 75.41 65.70 74.34
↓50 - 50 90.92 90.88 90.80 91.38 91.38 91.39 83.16 78.40 88.21
25 25 92.69 92.69 92.73 90.89 90.88 90.92 81.35 76.30 85.80
50 - 88.80 88.80 88.85 90.56 90.56 90.55 77.22 71.40 80.75
↓25 - 25 93.58 93.58 93.64 92.03 92.02 92.06 82.83 79.30 88.31
25 - 92.01 92.01 92.03 92.84 92.85 92.85 82.50 79.10 87.65
- - - 92.49 92.47 92.42 91.54 91.53 91.51 85.64 81.90 89.40
Table 6. Experimental results of fine-tuned LLaVA-v1.5-13B
Missing Rate (%) Modality Weibo Weibo21 Pheme
Text Image ACC M-F1 AUC ACC M-F1 AUC ACC M-F1 AUC
↓100 - 100 80.41 79.87 79.66 86.18 86.29 85.95 53.47 53.44 66.50
25 75 76.72 76.08 75.50 81.14 80.68 81.27 48.35 48.06 62.92
50 50 70.44 69.71 68.43 76.75 76.01 76.90 47.03 46.68 61.65
75 25 67.99 67.40 66.60 68.94 68.03 69.09 43.56 43.08 58.01
100 - 54.13 55.33 47.43 55.45 49.75 55.72 35.15 33.85 50.22
↓75 - 75 81.71 81.25 81.17 86.67 86.47 86.67 52.48 52.43 65.64
25 50 77.34 76.76 76.35 81.30 80.83 81.43 50.17 50.03 64.02
50 25 70.99 70.26 69.04 77.72 77.09 77.86 47.36 47.09 61.53
75 - 68.12 67.65 67.17 69.92 69.17 70.05 47.19 47.15 59.19
↓50 - 50 82.59 82.21 82.21 86.99 86.81 87.09 54.46 54.45 66.85
25 25 77.88 76.99 77.33 81.46 80.99 81.60 51.65 51.58 65.06
50 - 71.54 70.03 70.89 77.40 76.76 77.54 48.84 48.74 62.06
↓25 - 25 83.00 82.62 82.61 86.99 86.81 87.09 55.12 55.12 67.15
25 - 77.68 76.85 77.15 81.46 80.99 81.60 52.81 52.77 65.87
- - - 83.00 82.67 82.64 86.50 86.29 86.61 57.10 57.07 68.53
Table 7. Experimental results of fine-tuned Qwen-VL-7B
Missing Rate (%) Modality Weibo Weibo21 Pheme
Text Image ACC M-F1 AUC ACC M-F1 AUC ACC M-F1 AUC
↓100 - 100 77.88 76.60 77.21 77.07 75.93 77.26 50.17 49.99 64.36
25 75 72.90 70.60 72.08 67.32 63.87 67.57 45.87 45.37 61.01
50 50 68.67 65.31 67.74 62.76 57.52 63.05 46.04 45.62 60.78
75 25 65.46 62.32 64.60 57.40 49.97 57.72 42.08 41.45 56.80
100 - 53.65 49.24 54.65 49.76 33.51 50.16 36.80 36.51 48.46
↓75 - 75 79.32 78.40 78.74 77.40 76.31 77.58 49.67 49.48 63.84
25 50 74.40 72.54 73.64 68.78 65.67 69.03 48.51 48.27 62.86
50 25 68.94 65.77 68.04 64.23 59.56 64.51 47.03 46.74 61.30
75 - 66.69 64.48 65.96 56.42 48.52 56.75 45.38 45.27 57.74
↓50 - 50 80.48 79.71 79.94 77.56 76.49 77.74 52.15 52.10 65.23
25 25 74.33 72.47 73.58 69.43 66.60 69.43 50.00 49.87 63.73
50 - 70.65 68.25 69.84 63.09 57.90 63.38 48.35 48.20 61.89
↓25 - 25 80.89 80.20 80.37 77.40 76.31 77.58 53.14 53.12 65.93
25 - 75.09 73.52 74.38 68.62 65.53 68.87 51.32 51.26 64.48
- - - 81.71 81.12 81.22 76.91 75.74 77.09 55.12 55.12 67.15

### v2

@@ 4.6 Ablation Study
1) To validate the effectiveness and robustness of the modality-specfic multi-expert reasoning paradigm in MMLNet for the MFDN task, we conducted ablation studies by sequentially removing the textual expert ${L}_{c}^{h}$, the visual expert ${L}_{c}^{r}$, and the multimodal expert ${L}_{c}^{f}$. The framework achieved the best performance when all experts collaborated, demonstrating that all modal experts play indispensable roles in discerning the authenticity of news. Notably, removing the ${L}_{c}^{f}$ resulted in a significant performance drop, which highlights the critical importance of modeling the distribution of cross-modal interaction features for the success of MMR.
@@
4.7 Analysis on Modality Missing Learning.
Answear to RQ4. As shown in Figure 7(a), our modality missing learning is more suitable for MMR tasks with incomplete modalities. On the Weibo dataset with 50% text and 50% image modality missing, we conducted the following ablation experiments on modality missing learning:
@@
  • w/o Weight. MMLNet with modality missing learning but without the weight in Eq18.
@@
We visualized the multimodal features before classification via T-SNE [^5] https://github.com/mxl1990/tsne-pytorch in Figure 7. The ‘w/ MML’ line shows that MMLNet with modality missing learning achieves up to an 4% higher accuracy than without it, indicating our modality missing learning is beneficial in modality missing scenarios. ‘ w/o Weight’ is second-best, showing that positive and negative samples based on news authenticity labels improve the MMLNet model’s ability to detect misinformation. ‘w/o MML, w/ MCL’ has similar performance to ‘w/o MML’, but Figure 7(d) shows that vanilla multimodal contrastive learning has the worst visualization results, as modality missing undermines vanilla multimodal contrastive learning. This occurs because when image and text modality information is missing, the vanilla multimodal contrastive learning is misled by the missing information and cannot effectively distinguish between positive and negative samples.
4.8 Case Study
@@
Answear to RQ5. In Figure 8, we provide a case study on MMR to provide a detailed analysis with some interpretability. Analyzing the left 4 columns, although most baseline models are able to distinguish misinformation when the modality information is complete, they cannot make accurate judgments when the modality information is incomplete. However, MMLNet effectively corrects the impact of missing modality information on model performance by combining distributions from different modality experts. For example, when text modality absence predominates, the lack of text modality information causes erroneous judgments by the text expert, but the image modality expert and multimodal expert distributions compensate for this, improving MMLNet’s robustness in MMR.
@@
Answear to RQ6. Although MMLNet has achieved significant improvements in the incomplete modality MMR task, it still has limitations. When modality information is severely missing, it cannot effectively distinguish the authenticity of news. As shown in the sixth column of Figure 8, when the text modallity information is missing to a large extent, the model will mistakenly judge the news as true, mainly because of the loss of key person information. Similarly, when the image modal information is missing to a large extent, effective information cannot be obtained from the image modality. As shown in the eighth column of Figure 8, when the text information on the image is lost, resulting in semantic inconsistency between the image modality and the text modality, the model mistakenly judges the sample as fake news.
5 Conclusion and Discussion
@@
In this paper, we focus on the impact of modality missing information on the robustness of multimodal misinformation recognition. These issues are common in practical applications, yet they have been insufficiently explored in previous studies. We propose MMLNet, which models the distribution of different modalities through multi-expert learning. The modality missing adapter compensates for missing modality information, while incomplete modality learning enhances the robustness of feature distributions. Multi-expert collaboration enables accurate identification of misinformation. Extensive experiments and ablation studies under both complete and incomplete modalities demonstrate the effectiveness of our approach. Our future work will explore more missing modality scenarios in MMR tasks, such as video and audio modalities, to further improve the adaptability of our method. Meanwhile, we will study how to design more modality missingness to better fit real-world contexts.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No. 62376121 and No. 62501644.
Declarations
  • Conflict of interest. The authors declare that they have no Conflict of interest.
  • Data availability. The datasets used in this study are publicly available from the following sources: the Weibo dataset link, the Weibo21 dataset link, and the Pheme dataset link.
  • Code availability.
    https://github.com/zhyhome/MMLNet